OpenStudy (anonymous):

MEDALS!!!!!!!!!!!!!! Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - 2y = 6 2x + y = 5 -4x - 2y = 10 -4y = 4 -4x - 5y = -1 -7y = 5 3x + 2y = 6 7x = 12 4x + 2y = 10 8x = 16

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

the solution to the first system is x=2, y=1. Substituting these values in the other equations yields the answer

OpenStudy (anonymous):

thanks so much!!!! could you help me with some more questions?

OpenStudy (anonymous):

I can try

OpenStudy (anonymous):

ok great let me copy and paste them real quick

OpenStudy (anonymous):

Solve the system of equations. 5x + y = 9 3x + 2y = 4 (-2, 5) (1, 4) (2, -1) (4, -4)

OpenStudy (anonymous):

(2,-1)

OpenStudy (anonymous):

ok next one :D Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (anonymous):

(1,3)

OpenStudy (anonymous):

ok then we have this one :) Choose the equivalent system of linear equations that will produce the same solution as the one given below. 4x - y = -11 2x + 3y = 5 -4x - 9y = -19 -10y = -30 4x + 3y = 5 2y = -6 7x - 3y = -11 9x = -6 12x - 3y = -33 14x = -28

OpenStudy (anonymous):

we could come back to that one later, how about this one? A gym offers regular memberships for $80 per month and off-peak memberships for $60 per month. Last month, the gym sold a total of 420 memberships for a total of $31,100. How many memberships sold were regular memberships? 125 140 235 295

OpenStudy (anonymous):

a and d seem to be correct for the previous one

OpenStudy (anonymous):

ok thank you, so its either a or d

OpenStudy (anonymous):

and how bout the gym one?

OpenStudy (anonymous):

295

OpenStudy (anonymous):

ok r u up for a couple more???

OpenStudy (anonymous):

hmm give me a sec. need to shake off my brain muscles lol

OpenStudy (anonymous):

hahaha ok lol, no need to hurry :) Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -2 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

3rd option

OpenStudy (anonymous):

ok great, now i have two last questions then im out of your hair

OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 14 x + 3y = 9 (1, 12) (-1, -12) (12, -1) (12, 1)

OpenStudy (anonymous):

And also this one : Determine which system below will produce infinitely many solutions. 2x + 5y = 24 2x + 5y = 42 3x - 2y = 15 6x + 5y = 11 4x - 3y = 9 -8x + 6y = -18 5x - 3y = 16 -2x + 3y = -7

OpenStudy (anonymous):

(12,-1)

OpenStudy (anonymous):

3rd option

OpenStudy (anonymous):

omygosh, i just submitted it and i got a 100%!!!!!! :D Yay thank you so much!!!!!!! :DDDDDD

OpenStudy (anonymous):

You are the best person ever!!!!!!!!! :D